# Batch Predictive Inference

### v4

@@ Abstract
Constructing prediction sets with coverage guarantees for unobserved outcomes is a core problem in modern statistics. Methods for predictive inference have been developed for a wide range of settings, but usually only consider test data points one at a time. Here we study the problem of distribution-free predictive inference for a batch of multiple test points, aiming to construct prediction sets for functions—such as the mean or median—of any number of unobserved test datapoints. This setting includes constructing simultaneous prediction sets with a high probability of coverage, and selecting datapoints satisfying a specified condition while controlling the number of false claims.
For the general task of predictive inference on a function of a batch of test points, we introduce a methodology called batch predictive inference (batch PI), and provide a distribution-free coverage guarantee under exchangeability of the calibration and test data. Batch PI requires the quantiles of a rank ordering function defined on certain subsets of ranks. While computing these quantiles is NP-hard in general, we show that it can be done efficiently in many cases of interest, most notably for batch score functions with a compositional structure—which includes examples of interest such as the mean—via a dynamic programming algorithm that we develop. Batch PI has advantages over naive approaches (such as partitioning the calibration data or directly extending conformal prediction) in many settings, as it can deliver informative prediction sets even using small calibration sample sizes. We illustrate that our procedures provide informative inference across the use cases mentioned above, through experiments on both simulated data and a drug-target interaction dataset.
@@ 1 Introduction
Consider a supervised learning setting where we have a dataset $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ drawn from $P_{X,Y}=P_{X}\times P_{Y\mid X}$ on $\mathcal{X}\times\mathcal{Y}$ and a batch of new test inputs $X_{n+1},\ldots,X_{n+m}$ from $P_{X}$. Our task is to predict and make inference for the unobserved outcomes $Y_{n+1},\ldots,Y_{n+m}$, or more generally learn about the conditional distributions $P_{Y_{n+1}\mid X_{n+1}},\ldots,P_{Y_{n+m}\mid X_{n+m}}$. This setting includes both regression and classification. Beyond point predictions, it is of significant interest to construct prediction sets for various functions of the unobserved outcomes $Y_{n+1},\ldots,Y_{n+m}$. For example, given a regression function $\hat{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$ trained using a subset of the data, one might aim to construct a prediction set of the form $\widehat{C}_{n}(X_{n+1})=[\hat{\mu}(X_{n+1})-A,\hat{\mu}(X_{n+1})+A]$ for some $A>0$, which satisfies the marginal coverage guarantee $\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\}\geqslant 1-\alpha$, for a predetermined level $\alpha\in(0,1)$.
Distribution-free inference aims to achieve such inferential targets without imposing distributional assumptions on the sampling distribution $P_{X,Y}$, and dates back at least to the pioneering works of Wilks (1941), Wald (1943), Scheffe and Tukey (1945), and Tukey (1947, 1948). For example, conformal prediction (e.g., Saunders et al., 1999; Vovk et al., 1999, 2005, etc.) provides a general framework for achieving marginal coverage under exchangeability. Many recent works have explored the possibility of improving or generalizing this framework to achieve stronger targets, reduce computational costs, or enable inference with non-exchangeable data, etc, see Section 1.3. However, method development for joint inference on functions of multiple test points has been limited.
In this work, we develop methodology for distribution-free joint inference on multiple test points. This is a general problem, and includes as special case several examples of interest:
@@
    Prediction sets for multiple unobserved outcomes. Consider constructing an algorithm $\widehat{C}_{n}$ that likely obtains at least $1-\delta$ empirical coverage over the test set, i.e., $\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\delta}\right\}\geqslant 1-\alpha$. Compared to applying conformal prediction separately to individual test points to obtain $\mathbb{P}\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\}\geqslant 1-\alpha^{\prime}$, $j=1,\ldots,m$, for some $\alpha^{\prime}$, the above coverage guarantee directly states that most prediction sets cover the true outcome.
@@
    Inference on the mean. Consider predicting the mean of the test outcomes via a prediction set $\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})$ such that $\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}Y_{n+j}\in\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})}\right\}\geqslant 1-\alpha$. This problem had a range of use cases and we illustrate it in a problem of inference on the mean of counterfactual variables. When the size $m$ of the test data set is small, methods based on concentration inequalities can generally be conservative, producing wide intervals, in contrast, as we demonstrate empirically, our methods can still be informative.
@@
    Selection of of datapoints with error control. Consider selecting test datapoints in the test set whose responses satisfy a specific condition, such as $Y_{n+j}>c$ for a predetermined threshold $c$. As $(Y_{n+j})_{1\leqslant j\leqslant m}$ are unobserved, a potential approach is to construct a selection criterion based on the training and calibration data, e.g., of the form $\hat{\mu}(X_{n+j})>\hat{T}$. One possible inferential target is the control of the probability of making more than $k$ errors at level $\alpha$, i.e., $\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})>\hat{T},Y_{n+j}\leqslant c}\right\}>k}\right\}<\alpha$, where $k$ is a predetermined target error bound. This is analogous to the notion of family-wise error rate (FWER) control in multiple hypothesis testing.
We provide more details on the above examples in Section 3. The examples turn out to be special cases of the following general problem: Given the calibration data $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{1\leqslant i\leqslant n}$ and a function $g$ that takes the (multi)set of test observations as the input, construct a prediction set $\widehat{C}(\mathcal{D}_{n})$ that satisfies
@@
Notation. We write $\mathbb{R}$ to denote the set of real numbers and $\mathbb{R}_{\geqslant 0}$ to denote the set of nonnegative reals. The set of positive integers is denoted by $\mathbb{N}$. For a positive integer $n\in\mathbb{N}$, we write $[n]$ to denote the set $\{1,2,\ldots,n\}$ and for any $a,b\in[n]$ with $a\leqslant b$ write $X_{a:b}$ to denote the vector $(X_{a},X_{2},\ldots,X_{b})^{\top}$. We will denote the all ones vector of size $m$ as $1_{m}$. For a function $f:A\rightarrow B$, We write $\mathrm{Im}(f)$ to denote the image of a function $f$, and $f\big{|}_{C}$ to denote the restriction of $f$ to $C\subset A$. For a real number $x$, we write $\lfloor x\rfloor$, $\lceil x\rceil$, and $\text{round}(x)$ to denote the floor, ceiling, and rounding of $x$ (with 1/2 rounding up) to the nearest integer, respectively. We let $a_{+}=\max\{a,0\}$ for a real number $a\in\mathbb{R}$. We denote the number of ways to choose $r$ items with replacement from $n$ items as ${}_{n}\mathrm{H}_{r}$.
We write $\sum_{i=1}^{k}p_{i}\delta_{v_{i}}$ to denote the discrete distribution with support $\{v_{1},v_{2},\ldots,v_{k}\}$ and the probability masses $(p_{1},p_{2},\ldots,p_{k})$. For a distribution $P$, we define $Q_{\tau}(P)=\inf\{t\in\mathbb{R}:\mathbb{P}_{{X\sim Q}}\left\{{X\leqslant t}\right\}\geqslant\tau\}$ and $Q_{\tau}^{\prime}(P)=\sup\{t\in\mathbb{R}:\mathbb{P}_{{X\sim P}}\left\{{X\leqslant t}\right\}\leqslant\tau\}$. For an event $E$, we write ${\mathbbm{1}}\left\{{E}\right\}$ to denote its corresponding indicator variable. All objects (sets and functions) considered will be measurable with respect to appropriate sigma-algebras (typically the Borel sigma-algebra generated by open sets), which will not be mentioned further. For a set $D$, $\mathcal{P}(D)$ denotes its power set; or the Borel sigma algebra on $D$ if that is well-defined. We write $\mathcal{N}(\mu,\sigma^{2};[a,b])$ to denote the truncated normal distribution with mean $\mu$, variance $\sigma^{2}$, and truncation set $[a,b]$.
@@ 1.1 Main contributions
    Batch predictive inference (batch PI): We develop the batch predictive inference (batch PI): methodology for distribution-free inference on a function of multiple unobserved test outcomes. Our targets include a broad range of functions satisfying a certain monotonicity property, such as the mean or quantiles. Furthermore, we extend this approach to achieve simultaneous inference on multiple quantiles of test scores. Batch PI can provide useful inference when the calibration dataset size is comparable to—or even smaller than—the test size, a scenario in which we show that naive approaches fail.
@@
    Efficient algorithms for the batch PI procedure: We show that the batch PI procedure is generally NP-hard to compute, but it can be simplified for many target functions of practical interest, such as the mean and quantiles. For quantiles, and more generally for “sparse” functions depending only on a few quantiles, we establish how the computational burden can be reduced substantially, making the approach feasible in routine applications. For the mean, and more generally for functions satisfying a certain compositional structure, we present a polynomial-time dynamic programming algorithm for batch PI.
@@ 1.2 Problem setting
We observe data points $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$, with $(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}$ for $i\in[n]$, where $\mathcal{X}$ is a feature space and $\mathcal{Y}$ is an outcome space. Here and below, sets refer to multisets, and allow repetitions of elements. We view $\mathcal{D}_{n}$ as a calibration dataset, in the sense that we assume access to another independent dataset where one can build a predictor $\hat{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$. The calibration dataset $\mathcal{D}_{n}$ is used to construct a prediction set, leveraging $\hat{\mu}$. The prediction set is applied to the test dataset, which consists of $m\geqslant 1$ observed test features $X_{n+1},X_{n+2},\ldots,X_{n+m}$ without their corresponding outcomes $Y_{n+1},Y_{n+2},\ldots,Y_{n+m}$. We denote each data point as $Z_{i}=(X_{i},Y_{i})$, for $i\in[n+m]$.
Given a function $g:\{Z_{n+1},\ldots,Z_{n+m}\}\mapsto g(\{Z_{n+1},\ldots,Z_{n+m}\})$ of interest of the set of test points $\{Z_{n+1},\ldots,Z_{n+m}\}$, our goal is to construct a prediction set for the unobserved value $g(\{Z_{n+1},\ldots,Z_{n+m}\})$; which depends on the unobserved outcomes $Y_{n+1},Y_{n+2},\ldots,Y_{n+m}$. Specifically, we aim to construct a procedure $\widehat{C}:(\mathcal{X}\times\mathcal{Y})^{n}\rightarrow\mathcal{P}(\mathbb{R})$ such that
@@
holds for a predefined level $\alpha\in(0,1)$. We are interested in a general setting where $m$ is not necessarily significantly smaller than the calibration set size $n$ (in contrast to cases with trivial solutions, as we will describe later), and may even be larger. We will argue that this setting covers a wide range of important scenarios.
We now need some notations: For any vector $v\in\mathbb{R}^{m}$, let $v_{\uparrow}=(v_{(1)},\ldots,v_{(m)})$ be the vector $v$ sorted in a non-decreasing order, so that $v_{(1)}\leqslant\ldots\leqslant v_{(m)}$. For $z\in(\mathcal{X}\times\mathcal{Y})^{m}$ and a “score” function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, define $s(z)=(s(z_{1}),s(z_{2}),\ldots,s(z_{m}))$ by element-wise application of $s$. We denote $S_{i}=s(Z_{i})$ for all $i\in[m+n]$. Let $\mathbb{R}^{m}_{\uparrow}=\{x\in\mathbb{R}^{m}:x_{1}\leqslant x_{2}\leqslant\ldots\leqslant x_{m}\}$ be the set of monotone non-increasing vectors. For two vectors $u=(u_{1},\ldots,u_{d})^{\top},v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$, we write $u\preceq v$ if $u_{i}\leqslant v_{i}$ for all $i=1,2,\ldots,d$.
@@
There is a batch score function [^2] Let $\inf s=\inf\{s(x,y):(x,y)\in\mathcal{X}\times\mathcal{Y}\}$ and $\sup s=\sup\{s(x,y):(x,y)\in\mathcal{X}\times\mathcal{Y}\}$. When $s$ is unbounded, we will also need the function $h$ to be defined for all values $s_{1}\leqslant\ldots\leqslant s_{m}$ such that $s_{i}\in(\inf s,\sup s)$ for all $i\in\{2,\ldots,m-1\}$ and either $s_{1}=\inf s$ or $s_{m}=\sup s$. We will define $h(-\infty,s_{2},\ldots,s_{m})=-\infty$ if $s_{1}=\inf s=-\infty$, and $h(s_{1},\ldots,s_{m-1},\infty)=\infty$ if $s_{m}=\sup s=\infty$. $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$ and a (non-batch, per-datapoint) score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, such that
@@
Condition 1 covers a broad range of targets, from the mean $h(s_{1},\ldots,s_{m})=\frac{s_{1}+\ldots+s_{m}}{m}$ and the $q$-th quantile $h(s_{1},\ldots,s_{m})=s_{(\lceil(qm)\rceil)}$, $q\in(0,1)$, to more general targets such as the truncated mean or the proportion of scores exceeding a certain threshold. In many settings, $h$ represents a fixed function of interest, while $s$ is typically constructed using a separate dataset. For instance, in regression tasks, we can consider nonconformity scores such as $s:(x,y)\mapsto|y-\hat{\mu}(x)|$, where $\hat{\mu}$ is fitted on a separate dataset.
If the cardinality of $\mathcal{X}\times\mathcal{Y}$ is at most that of $\mathbb{R}$ (which holds for instance in the typical case that $\mathcal{X}\subset\mathbb{R}^{d}$ for some positive integer $d\geqslant 0$, and $\mathcal{Y}$ is either discrete or $\mathbb{R}$), then there is an injective map $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. Let $\mathcal{S}\subset\mathbb{R}$ be the image of $s$. Then, for any $v\in\mathbb{R}^{m}_{\uparrow}\cap\mathcal{S}^{m}$, we can define $h(v)=g(\{s^{-1}(v_{1}),\ldots,s^{-1}(v_{m})\})$. For $v\in\mathbb{R}^{m}_{\uparrow}\setminus\mathcal{S}^{m}$, we can define $h(v)$ arbitrarily. Since $s$ is injective, $h$ is well-defined. Moreover, (2) holds by definition. Thus, if the cardinality of $\mathcal{X}\times\mathcal{Y}$ is at most that of $\mathbb{R}$, (2) holds, and only the monotonicity property (3) imposes a condition.

### v5

@@ Abstract
Constructing prediction sets with coverage guarantees for unobserved outcomes is a core problem in modern statistics. Methods for predictive inference have been developed for a wide range of settings, but usually only consider test data points one at a time. Here we study the problem of distribution-free predictive inference for a functions of batch of multiple test points, aiming to construct prediction sets for functions—such as the mean or median—of any number of unobserved test datapoints. This setting includes constructing simultaneous prediction sets with a high probability of coverage, and selecting datapoints satisfying a specified condition (e.g., being large) while controlling the number of false claims. Here, for the general task of predictive inference on a function of a batch of test points, we introduce a methodology called batch predictive inference (batch PI), and provide a distribution-free coverage guarantee under exchangeability of the calibration and test data. Batch PI requires the quantiles of a rank ordering function defined on certain subsets of ranks. While computing these quantiles is NP-hard in general, we show that it can be done efficiently in many cases of interest, most notably for batch score functions with a compositional structure—which includes examples of interest such as the mean—via a dynamic programming algorithm that we develop. Batch PI has advantages over baseline approaches (such as partitioning the calibration data or directly extending conformal prediction) in many settings, as it can deliver informative prediction sets even using small calibration sample sizes. We illustrate that our procedures provide informative inference across the use cases mentioned above, through experiments on both simulated data and a drug-target interaction dataset.
@@ 1 Introduction
Consider a supervised learning setting where we have a dataset $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ drawn from $P_{X,Y}=P_{X}\times P_{Y\mid X}$ on $\mathcal{X}\times\mathcal{Y}$ and a batch of new test inputs $X_{n+1},\ldots,X_{n+m}$ from $P_{X}$. Our task is to predict and make inference for the unobserved outcomes $Y_{n+1},\ldots,Y_{n+m}$. This setting includes both regression and classification. Beyond point predictions, it is of significant interest to construct prediction sets for various functions of the unobserved outcomes $Y_{n+1},\ldots,Y_{n+m}$. For example, given a regression function $\hat{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$ trained using a subset of the data, one might aim to construct a prediction set for $Y_{n+1}$ of the form $\widehat{C}_{n}(X_{n+1})=\hat{\mu}(X_{n+1})\pm(\text{constant})$, which satisfies the marginal coverage guarantee $\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\}\geqslant 1-\alpha$, for a predetermined level $\alpha\in(0,1)$.
Distribution-free inference aims to achieve such inferential targets without imposing distributional assumptions on the sampling distribution $P_{X,Y}$, and dates back at least to the pioneering works of Wilks (1941), Wald (1943), Scheffe and Tukey (1945) in the 1940s, and Tukey (1947, 1948). For example, conformal prediction (e.g., Saunders et al., 1999; Vovk et al., 1999, 2005, etc.) provides a general framework for achieving marginal coverage under exchangeability. Many recent works have explored the possibility of improving or generalizing this framework to achieve stronger targets, reduce computational costs, or enable inference with non-exchangeable data, etc, see Section 1.3. However, method development for joint inference on functions of multiple test points has been limited.
In this work, we develop methodology for distribution-free joint inference on multiple test points. At a high level, this problem is connected to two major areas of statistical research:
@@
    Simultaneous inference on multiple quantities. In many real-world problems, there are multiple quantities of interest for inference—e.g., multiple applicants for a job (Cohen et al., 2020; Barigozzi and Burani, 2016), patients undergoing screening or a particular treatment (Nielsen and Lang, 1999; Colombo, 2007), drug candidates in high-throughput screening (Mayr and Bojanic, 2009; Macarron et al., 2011), multiple endpoints in medical trials (Budig et al., 2024), weather or other variables in weather forecasting (Neeven and Smirnov, 2018; Messoudi et al., 2022; Sampson and Chan, 2024). For testing problems, a series of methods have been developed for multiplicity adjustment to obtain valid multiple testing procedures (see e.g., Lehmann and Romano, 2005b; Miller, 2012, etc). However, for predictive inference problems, methodological development remains limited. We will show that existing approaches often struggle to provide useful valid inferential guarantees in this setting.
@@
    Inference on a finite population. In applications such as survey studies and randomized trials (Kalton, 2020; Hariton and Locascio, 2018), researchers are often interested in analyzing a finite population rather than a hypothetical infinite population (see e.g., Abadie et al., 2020, etc)—for example, the distribution of treatment effects across the group of individuals who received the treatment. A school administrator may want to anticipate the effect of a new teaching method specifically on the students in a program, rather than on a hypothetical broader student population, see e.g., Kautz et al. (2017). Similarly, in the analysis of network data, researchers are often interested in understanding how a message or intervention spreads through a specific social network Newman (2018), and network models that include exchangeable feature observations have been studied (Mao et al., 2021).
More specifically, this problem includes several inferential tasks as special cases:
  1. 1.
    Inference on the mean of a test dataset; including on counterfactuals. Consider predicting the mean of the test outcomes via a prediction set $\widehat{C}_{n}$ such that
    $\mathbb{P}\Big{\{}\frac{1}{m}\sum\nolimits_{j=1}^{m}Y_{n+j}\in\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})\Big{\}}\geqslant 1-\alpha.$
    This problem has a range of use cases and we illustrate it in a problem of inference on the mean difference between counterfactual outcomes under different treatments. Specifically, consider a randomized trial where $A\in\{0,1\}$ denotes the treatment assignment, and $Y^{a=0}$ and $Y^{a=1}$ represent the counterfactual outcomes under control and treatment, respectively. For each individual $i=1,\ldots,n$, we observe the triplet $(X_{i},A_{i},(1-A_{i})Y_{i}^{a=0}+A_{i}Y_{i}^{a=1})$—that is, we observe only the counterfactual corresponding to the assigned treatment. Our goal is to construct a prediction set $\widehat{C}(\mathcal{D})$ for the mean of the unobserved counterfactual outcomes among the treated subgroup:
    $\mathbb{P}\Big{\{}\frac{1}{N^{1}}\sum_{i:A_{i}=1}Y_{i}^{a=0}\in\widehat{C}(\mathcal{D})\Big{\}}\geqslant 1-\alpha,$
    where $N^{1}=|\{i:A_{i}=1\}|$, and $\mathcal{D}$ here denotes the full set of the observed data. When the number of test datapoints is small, methods based on concentration inequalities can generally be conservative for the above problems, producing wide intervals. In contrast, as we demonstrate empirically, our methods can still be informative. It is also worth mentioning that our method also works for the median and other quantiles; and in particular for the median counterfactual.
  2. 2.
    Prediction sets for multiple unobserved outcomes. Consider constructing an algorithm $\widehat{C}_{n}$ that likely obtains at least $1-\delta$ empirical coverage over the test set, i.e.,
    $\mathbb{P}\Big{\{}\frac{1}{m}\sum\nolimits_{j=1}^{m}{\mathbbm{1}}\left\{{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})}\right\}\geqslant 1-\delta\Big{\}}\geqslant 1-\alpha.$
    Compared to applying conformal prediction separately to individual test points to obtain $\mathbb{P}\{Y_{n+j}\in\widehat{C}_{n}(X_{n+j})\}\geqslant 1-\alpha^{\prime}$, $j=1,\ldots,m$, for some $\alpha^{\prime}$, the above coverage guarantee directly states that most prediction sets cover the true outcome with a well-calibrated and pre-specified high-probability. For instance, this allows us to construct prediction sets for a machine learning classifier, such that for a test set of interest, most labels are covered with a given probability.
@@
    Selection of datapoints with error control. Consider selecting test datapoints in the test set whose responses satisfy a specific condition, such as $Y_{n+j}>c$ for a predetermined threshold $c$. As $(Y_{n+j})_{1\leqslant j\leqslant m}$ are unobserved, a potential approach is to construct a selection criterion based on the training and calibration data, e.g., of the form $\hat{\mu}(X_{n+j})>\hat{T}$. One possible inferential target is the control of the probability of making more than $k$ errors at level $\alpha$, i.e.,
    $\mathbb{P}\Big{\{}\sum\nolimits_{j=1}^{m}{\mathbbm{1}}\left\{{\hat{\mu}(X_{n+j})>\hat{T},Y_{n+j}\leqslant c}\right\}>k\Big{\}}<\alpha,$
    where $k$ is a predetermined target error bound. This is analogous to the notion of family-wise error rate (FWER) control in multiple hypothesis testing. As an example, we use this method to select promising drug-target pairs.
We provide more details on the above examples in Section 3. The examples turn out to be special cases of the following general problem: Given the calibration data $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{1\leqslant i\leqslant n}$ and a function $g:\mathcal{P}(\mathcal{X}\times\mathcal{Y})\rightarrow\mathbb{R}$ [^2] For a set $A$, we write $\mathcal{P}(A)$ to denote its power set. that takes the set of test observations as the input, construct a prediction set $\widehat{C}(\mathcal{D}_{n})$ that satisfies
@@
Notation. We write $\mathbb{R}$ to denote the set of real numbers and $\mathbb{R}_{\geqslant 0}$ to denote the set of nonnegative reals. The set of positive integers is denoted by $\mathbb{N}$. For a positive integer $n\in\mathbb{N}$, we write $[n]$ to denote the set $\{1,2,\ldots,n\}$ and for any $a,b\in[n]$ with $a\leqslant b$ write $X_{a:b}$ to denote the vector $(X_{a},X_{2},\ldots,X_{b})^{\top}$. We will denote the all ones vector of size $m$ as $1_{m}$. For a function $f:A\rightarrow B$, We write $\mathrm{Im}(f)$ to denote the image of a function $f$, and $f\big{|}_{C}$ to denote the restriction of $f$ to $C\subset A$. For a real number $x$, we write $\lfloor x\rfloor$, $\lceil x\rceil$, and $\text{round}(x)$ to denote the floor, ceiling, and rounding of $x$ (with 1/2 rounding up) to the nearest integer, respectively. We let $a_{+}=\max\{a,0\}$ for a real number $a\in\mathbb{R}$. We denote the number of ways to choose $r$ items with replacement from $n$ items as ${}_{n}\mathrm{H}_{r}$. Let $\mathbb{R}^{m}_{\uparrow}=\{x\in\mathbb{R}^{m}:x_{1}\leqslant x_{2}\leqslant\ldots\leqslant x_{m}\}$ be the set of monotone non-increasing vectors. For two vectors $u=(u_{1},\ldots,u_{d})^{\top},v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$, we write $u\preceq v$ if $u_{i}\leqslant v_{i}$ for all $i=1,2,\ldots,d$.
We write $\sum_{i=1}^{k}p_{i}\delta_{v_{i}}$ to denote the discrete distribution with support $\{v_{1},v_{2},\ldots,v_{k}\}$ and the probability masses $(p_{1},p_{2},\ldots,p_{k})$. For a distribution $P$, we define two types of quantile functions $Q_{\tau}(P)=\inf\{t\in\mathbb{R}:\mathbb{P}_{{X\sim P}}\left\{{X\leqslant t}\right\}\geqslant\tau\}$ and $Q_{\tau}^{\prime}(P)=\sup\{t\in\mathbb{R}:\mathbb{P}_{{X\sim P}}\left\{{X\geqslant t}\right\}\geqslant 1-\tau\}$ [^3] It holds that $Q_{\tau}^{\prime}(P)=-Q_{1-\tau}(-P)$, where $-P$ denotes the distribution of $-X$ when $X\sim P$.. For an event $E$, we write ${\mathbbm{1}}\left\{{E}\right\}$ to denote its corresponding indicator variable. All objects (sets and functions) considered will be measurable with respect to appropriate sigma-algebras (typically the Borel sigma-algebra generated by open sets), which will not be mentioned further. For a set $D$, $\mathcal{P}(D)$ denotes its power set; or the Borel sigma algebra on $D$ if that is well-defined. We write $\mathcal{N}(\mu,\sigma^{2};[a,b])$ to denote the truncated normal distribution with mean $\mu$, variance $\sigma^{2}$, and truncation set $[a,b]$.
@@ 1.1 Main contributions
    Batch predictive inference (batch PI): We develop the batch predictive inference (batch PI): methodology for distribution-free inference on a function of multiple unobserved test outcomes. Our targets include a broad range of functions satisfying a certain monotonicity property, such as the mean or quantiles. Furthermore, we extend this approach to achieve simultaneous inference on multiple quantiles of test scores. Batch PI can provide useful inference when the calibration dataset size is comparable to—or even smaller than—the test size, a scenario in which we show that baseline approaches fail.
@@
    Efficient algorithms for the batch PI procedure: We show that the batch PI procedure is generally NP-hard to compute, but it can be simplified for many target functions of practical interest, such as the mean and quantiles. For quantiles, and more generally for “sparse” functions depending only on a few quantiles, we establish how the computational burden can be reduced substantially, making the approach feasible in routine applications. For functions satisfying a certain compositional structure (e.g., the mean), we present a polynomial-time dynamic programming algorithm for batch PI.
@@ 1.2 Problem setting
We observe data points $\mathcal{D}_{n}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}\subset\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is a feature space and $\mathcal{Y}$ is an outcome space. Here and below, sets refer to multisets, and allow repetitions of elements. We denote $\mathcal{D}_{n}$ as a calibration dataset, in the sense that it will be used for inference, e.g., to construct a prediction set. We then observe a test dataset consisting of $m\geqslant 1$ test features $X_{n+1},\ldots,X_{n+m}$, for which the corresponding outcomes $Y_{n+1},\ldots,Y_{n+m}$ are not observed. We denote each data point as $Z_{i}=(X_{i},Y_{i})$, for $i\in[n+m]$.
Given a real-valued function $g:\mathcal{P}(\mathcal{X}\times\mathcal{Y})\rightarrow\mathbb{R}$ of interest taking as input a subset of $\mathcal{X}\times\mathcal{Y}$, our goal is to construct a prediction set for the unobserved value $g(\{Z_{n+1},\ldots,Z_{n+m}\})$; which depends on the unobserved outcomes $Y_{n+1},\ldots,Y_{n+m}$. Specifically, we aim to construct a procedure $\widehat{C}:(\mathcal{X}\times\mathcal{Y})^{n}\rightarrow\mathcal{P}(\mathbb{R})$ such that
@@
holds for a predefined level $\alpha\in(0,1)$, regardless of the sampling distribution. We are interested in a general setting where $m$ is not necessarily significantly smaller than the calibration set size $n$ (in contrast to cases with trivial solutions, as we will describe later), and may even be larger. We will argue that this setting covers a wide range of important scenarios.
We now need some notations: For any vector $v\in\mathbb{R}^{m}$, let $v_{\uparrow}=(v_{(1)},\ldots,v_{(m)})$ be the vector $v$ sorted in a non-decreasing order. For $z=(z_{1},\ldots,z_{m})\in(\mathcal{X}\times\mathcal{Y})^{m}$ and a “score” function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, define $s(z)=(s(z_{1}),s(z_{2}),\ldots,s(z_{m}))$ by element-wise application of $s$. We denote $S_{i}=s(Z_{i})$ for all $i\in[m+n]$.
@@
There is a batch score function [^4] Let $\inf s=\inf\{s(x,y):(x,y)\in\mathcal{X}\times\mathcal{Y}\}$ and $\sup s=\sup\{s(x,y):(x,y)\in\mathcal{X}\times\mathcal{Y}\}$. When $s$ is unbounded, we need the function $h$ to be defined for all values $s_{1}\leqslant\ldots\leqslant s_{m}$ such that $s_{i}\in(\inf s,\sup s)$ for all $i\in\{2,\ldots,m-1\}$ and either $s_{1}=\inf s$ or $s_{m}=\sup s$. We define $h(-\infty,s_{2},\ldots,s_{m})=-\infty$ if $s_{1}=\inf s=-\infty$, and $h(s_{1},\ldots,s_{m-1},\infty)=\infty$ if $s_{m}=\sup s=\infty$. $h:\mathbb{R}^{m}_{\uparrow}\rightarrow\mathbb{R}$ and a (non-batch, per-datapoint) score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, such that
@@
Condition 1 covers a broad range of targets, from the mean $h(s_{1},\ldots,s_{m})=\frac{s_{1}+\ldots+s_{m}}{m}$ and the $q$-th quantile $h(s_{1},\ldots,s_{m})=s_{(\lceil(qm)\rceil)}$, $q\in(0,1)$, to more general targets such as the truncated mean or the proportion of scores exceeding a certain threshold. In many settings, $h$ represents a fixed function of interest, while $s$ is typically constructed using a separate dataset. For instance, in regression tasks, we can consider nonconformity scores such as $s:(x,y)\mapsto|y-\hat{\mu}(x)|$, where $\hat{\mu}$ is fitted on a separate dataset. As a simpler example, one can consider $s(y)=y$ and $m=2$, with $h(s_{1},s_{2})=\frac{s_{1}+s_{2}}{2}$, where the goal becomes inference on the average of two test outcomes, $(Y_{n+1}+Y_{n+2})/2$.
As a remark, if the cardinality of $\mathcal{X}\times\mathcal{Y}$ is at most that of $\mathbb{R}$—e.g., $\mathcal{X}\subset\mathbb{R}^{d}$ for some positive integer $d\geqslant 1$ and $\mathcal{Y}=\mathbb{R}$—then (2) holds, and only the monotonicity property (3) imposes a condition. [^5] To see that, observe that in this case, there is an injective map $f:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. Let $\mathcal{I}\subset\mathbb{R}$ be the image of $f$. Then, for any $v\in\mathbb{R}^{m}_{\uparrow}\cap\mathcal{I}^{m}$, we can define $h(v)=g(\{f^{-1}(v_{1}),\ldots,f^{-1}(v_{m})\})$, and for $v\in\mathbb{R}^{m}_{\uparrow}\setminus\mathcal{I}^{m}$, we can define $h(v)$ arbitrarily. Since $f$ is injective, $h$ is well-defined, satisfying (2) by definition.